A firm is considering making a one-time investment of $12.0 million, payable in full today. It is expected this would increase the firm's free cash flows by $300,000 in one year, increasing by 3% each year thereafter, forever. So in two years, for example, this investment will result in $309,000 more cash for the firm. What is the IRR of this investment? A firm is considering making a one-time investment of $12.0 million, payable in full today. It is expected this would increas
Answer #1

NPV=-12*10^6+300000/(r-3%)

At IRR: NPV=0
-12*10^6+300000/(IRR-3%)=0
=>IRR=3%+300000/(12*10^6)
=>IRR=5.50%